OpenStudy (anonymous):

Anyone? Im really not good at this stuff. Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). A. y = x + 7 B. y = -3x + 3 C. y = x + 3 D. y = -3x + 7

OpenStudy (anonymous):

plug those points into this formula: y-y1=m(x-x1) y-6=-3(x+1)

OpenStudy (anonymous):

Solve for y

OpenStudy (anonymous):

y= 3-3x ?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

First muliply -3 and x which is -3x Right?

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

ok and what is -3 times 1?